OpenStudy (anonymous):

Write the equation in Standard form: 8 - 4(x - y) = 2x + 6

jimthompson5910 (jim_thompson5910):

8 - 4(x - y) = 2x + 6 8 - 4x + 4y = 2x + 6 - 4x + 4y = 2x + 6 - 8 - 4x + 4y - 2x = 6 - 8 - 6x + 4y = -2 6x - 4y = 2